OpenStudy (anonymous):

Can someone please help me? Judy drove 20 miles from her house to a theater at an average rate of 50 miles per hour. Greg drove from his house to the theater in 1/3 of the time it took Judy to drive to the theater, and they arrived at the same time. If Judy left the house at 7:30 pm, when did Greg leave the house? The correct answer is 7:46pm but I can't figure it out why.

OpenStudy (anonymous):

What is the explanation of question 11 .At a certain fruit stand ,the price of an apple is twice the price of an orange.For which of the following compinations of apples and oranges is the total price equal to the total price of 20 oranges? Indicate all such combinations. The answer is 2apple and 16 orange/3 apple and 14 orange/6 apple and 6 orange 8

OpenStudy (ganpat):

Ok, SO for Judy : Speed = distance / time.. 50 = 20/ t Therefor t = 2/5 hrs.. For Greg ; time taken is 1/3 (Judy) = 1/3 (2/5) = 2/15 Therefore time taken for judy = 7.30 + 2/5 * 60 = 7:30 + 0: 24 = 7: 54\ And time at Greg left = 7.54 - 2/15 * 60 = 7: 54 - 0:08 = 7: 46.. Here is the solution...

OpenStudy (ganpat):

what u want ?? u have the answers !!